OpenStudy (anonymous):

The number of people in a town of 10,000 who have heard a rumor started by a small group of people is given by N(t) = 10,000/(5+1245e^-.97t) How long before 1,000 people have heard it?

OpenStudy (anonymous):

set that sucker equal to 1000 and solve for \(t\)

OpenStudy (anonymous):

I've gotten to 10 = 5+1245e^-.97t But don't know what to do next. Do I use natural log now?

OpenStudy (anonymous):

\[\frac{ 10,000}{5+1245e^{-.97t}}=1000 \]

OpenStudy (anonymous):

subtract 5 from both sides

OpenStudy (anonymous):

ok lets see what the first steps are before getting to the log part

OpenStudy (anonymous):

your goal is to write \[e^{-.97t}=\text{some number}\] Then use the log to solve for \(t\)

OpenStudy (anonymous):

So I'll have to divide 5 by 1245?

OpenStudy (anonymous):

so treat it like \[ 10 = 5+1245x\] sovle for \(x\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

course if it was me, i would use technology, but you can do that on a test i guess

OpenStudy (anonymous):

did you get to \[e^{-.97t}=\frac{1}{249}\] yet ?

OpenStudy (anonymous):

I'm using my calculator. I have \[\frac{ 1 }{ 249 }\] =\[e ^{-.97t}\]

OpenStudy (anonymous):

yup two more steps

OpenStudy (anonymous):

Now use natural log?

OpenStudy (anonymous):

yes now

OpenStudy (whpalmer4):

You can do another step to make writing easier :-) \[249 = e^{0.97t}\]

OpenStudy (anonymous):

I got -5.52. Would I set that equal to natural log -.97t? Then divide by -.97?

OpenStudy (anonymous):

I divided -5.52 by -.97 and got 5.7. So t = 5.7?

OpenStudy (whpalmer4):

Yes, t = 5.6881 is what i got.